sqrt of x+9-3 = sqrt of x-6
Answer by mukhopadhyay(490)

sqrt(x+9) - 3 = sqrt(x-6)
Square both sides:
(x+9) - 6sqrt(x+9) + 9 = (x-6)
=> 6sqrt(x+9) = (x+9)+9-(x-6)
=> 6sqrt(x+9) = 24
=> sqrt(x+9) = 4
=> (x+9) = 16 (Square both sides again)
=> x = 7